OpenStudy (anonymous):

I have the surface area of a cube(1014 cm2) and i need to know how to get the side lengths out of the surface area. can anyone help? need this for an exam review for tomorrow

OpenStudy (kinggeorge):

We have the surface area, not the volume.

OpenStudy (anonymous):

want to go over the steps with me that would sure help a lot

OpenStudy (kinggeorge):

To find the side lengths, first notice that a square has 6 sides, and each has the same area. Thus, we first have to divide \(1014/6=169\; \text{cm}^2\).That's the are of a single side of the cube.

OpenStudy (anonymous):

the surface area = 6 * the surface of a side the surface of a side is side lenght^2

OpenStudy (kinggeorge):

Now we have the area of a side of a square, from there, we just have to take the square root \[\sqrt{169}=13 \text{ cm}\]So your side length should be exactly 13cm.

OpenStudy (anonymous):

thank you so much thats big help,!

OpenStudy (kinggeorge):

You're welcome.